OpenStudy (rogabian98):

For which equation is x = 9 a solution? A. 3x – 7 = –34 B. x – 9 = 1 C. 2003-05-02-00-00_files/i0020000.jpg D. 4x + 5 = 41

OpenStudy (gracygirl):

What do you think?

OpenStudy (gracygirl):

Do you have an idea of which one?

OpenStudy (rogabian98):

no

OpenStudy (gracygirl):

Okay, so... D. 4x + 5 = 41 If you plug in 9... 4(9) + 5 = 41 36 + 5 = 41 Does that make sense?